f(x)=[x-1/2]+[x+1/2]+2[-x]
range of the function?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

f(x)=[x-1/2]+[x+1/2]+2[-x]

let x=I +f  

 where I is integer part of x and 0<= f<1

f(x) =[I+f-1/2]+[I+f+1/2]+2[-I-f]

        =I +[f-1/2] + I +[f+1/2]  -2I +2[-f]

f(x)=[f-1/2]+[f+1/2]+2[-f]

case 1    0=< I <1/2

f(x)= -1 +0 -2   =-3

case 2    1/2=< I <1

f(x)  = 0 +1 -2  =-1

so range is  -1 and -3


 

 

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin